# Mean-field dynamics of attractive resource interaction: From uniform to aggregated states

### v1

@@
Mean-field dynamics of attractive resource interaction: From uniform to aggregated states
We introduce and study a nonlinear discrete dynamical system describing the evolution of a resource distribution among interacting agents. The model generalizes several classical mean-field and opinion–dynamics frameworks and is defined on the standard simplex, where each coordinate evolves according to an interaction rule depending on pairwise preference functions. We provide a complete analytical description of the long-term behavior of the system. First, we establish monotonicity properties and show that the dynamics always remains in a positively invariant region determined by initial preferences. We prove the existence of a unique fixed point for any admissible parameter set and derive an explicit closed-form formula for the equilibrium in arbitrary dimension. We then analyze the local stability of the fixed point and identify parameter regimes leading to aggregation or uniform distributions. Finally, we characterize all possible asymptotic scenarios and show that, despite the nonlinear structure, the system does not exhibit oscillatory or chaotic behavior: every trajectory converges to the unique equilibrium. The results provide a full qualitative theory for this class of monotone resource–interaction models and offer a mathematical explanation for the transition from uniform to aggregated states.
Keywords: discrete dynamical systems, attractive interaction, stability analysis, transcritical bifurcation, resource distribution, mean-field models.
@@ 1 Introduction
The type of interaction between components – repulsive or attractive – critically affects the system’s dynamics. Models with repulsive interaction, which often describe competition, typically lead to the system’s tendency to avoid resource concentration. In contrast, models with attractive interaction, which are the focus of this paper, can describe processes of cooperation, gravitation towards centers of influence, or capital concentration. It is such systems that are the subject of our investigation. While the term ’conflict’ might seem counterintuitive in a model with attractive interaction, it represents a competition for dominance or influence within the state space $\Omega$. In this context, attraction towards more influential components is a key strategic element of the conflict, leading to the aggregation of resources rather than their uniform dispersion. This connects to the foundational concept of ’conflict’ between probability measures introduced by Volodymyr Koshmanenko [4, 5]. While the original work focused on quantifying divergence, the attractive dynamics explored here demonstrate how such interactions can lead to ’conflict resolution’ through convergence towards a single, stable equilibrium state. The mathematical modeling of interaction between conflicting systems, especially concerning the role of external support, is a subject of ongoing research [6].
@@
An alternative and powerful interpretation is as a system of interacting agents striving to reach consensus . In this view, the state vector represents a distribution of opinions, making the model an example of a discrete-time multi-agent system [10, 11]. Aggregated states then represent opinion polarization or fragmentation, phenomena widely studied in sociophysics [12, 13, 2].
Finally, the model shares properties with econophysics models of wealth distribution. The conservation law $\sum p_{i}^{t}=1$ mirrors a closed economy. While stochastic agent-based models like that of Chakraborti & Chakrabarti (2000) lead to Boltzmann-Gibbs distributions [14], our deterministic model provides an alternative mechanism for wealth aggregation, where the parameters $c_{i}$ can be interpreted as the ’economic attractiveness’ of an agent, analytically determining the final distribution.
Next, we construct a model of a dynamical system that describes the uniform or aggregated distribution of a certain resource in a given state space $\Omega=\{\omega_{1},\omega_{2},\dots,\omega_{n}\}$, $n>1$. We fix a certain discrete probability measure $\mu$, which corresponds to a specific distribution. The value of this measure,
$\mu(\omega_{i})=:p_{i}\geq 0,\quad i=\overline{1,n},$
is interpreted as the share of the resource or the relative weight of the i-th component of the system. By construction, $\sum\limits_{i}p_{i}=1$. Therefore, the sequence $\mu(\omega_{i})$ defines a stochastic vector in ${\mathbb{R}}_{+}^{n}$: ${\bf p}=(p_{1},p_{2},\dots,p_{n})$.
@@
It is easy to verify that these values also form a stochastic vector ${\bf{r}}=(r_{1},r_{2},\dots,r_{n})\in{\mathbb{R}_{+}^{n}}$. The quantities $\nu(\omega_{i})$ can be interpreted as an average distribution (mean field) of $\mu$ over the subset $\Omega\setminus\omega_{i}$. This mean-field approach is widely used in statistical physics to describe systems where individual components interact with an average property of the entire system rather than with each other individually [15].
@@
$p^{t+1}_{i}=\frac{p^{t}_{i}\big(1+{c_{i}}r^{t}_{i}\big)}{z^{t}},\quad 0<c_{i}\leq 1,\quad i=\overline{1,n},$ (2)
where
$r^{t}_{i}=\frac{1-p^{t}_{i}}{n-1},$ (3)
and the normalizing denominator $z^{t}$ ensures the stochasticity of the vector ${\bf{p}}^{t+1}$. The set of constants $c_{i}$ is interpreted as the time-invariant ($t$-independent) favorable conditions associated with each region $\omega_{i}$, respectively.
@@
It can be noted that the ’game for power’ model proposed in [16] is a prime example of a system with repulsion, where agents seek to minimize the influence of others. The model presented in this manuscript, featuring attractive interaction ((2)), can thus be interpreted as a ’game for cooperation’ or alliance formation, where an agent’s benefit increases by gravitating towards resources controlled by others. The choice of attractive interaction fundamentally changes the system’s dynamics and significantly affects its properties.
@@
2 Uniformly distributed limit states
Assume that in formulas (2), all $c_{i}=1$ for all $i=\overline{1,n}$, meaning favorable conditions are identical for each state $\omega_{i}$ and remain unchanged over time. Then equation (2) can be rewritten as
$p^{t+1}_{i}=\frac{p^{t}_{i}\big(1+r^{t}_{i}\big)}{1+\left({\bf{p}}^{t},{\bf{r}}^{t}\right)},\quad i=\overline{1,n}.$ (4)
@@ 2 Uniformly distributed limit states
The limit point where $p_{i}^{\infty}=\frac{1}{n}$ is known in the multi-agent systems literature as the consensus state [11, 18, 19]. It is a direct analogue to the outcome of classic linear consensus protocols, such as the DeGroot model, where iterative averaging on a fully connected graph leads to a system-wide average [20]. The model presented here, in this special case of uniform $c_{i}$, can thus be viewed as a novel, nonlinear consensus protocol.
@@
Every trajectory of the dynamic conflict system (1) with an arbitrary pair of initial stochastic vectors ${\bf{p}},{\bf{r}}\in{\mathbb{R}}_{+}^{n}$} $(n>1)$, whose evolution is given by formulas (5), converges to a limit state, which is a fixed point
@@ 2. $2)$
The limit state is stable only in the case ${\bf{p}}^{\infty}={\bf{r}}^{\infty}=\left(\frac{1}{n},\frac{1}{n},\dots,\frac{1}{n}\right)$.
@@
If $p_{i}\geq p_{j}$ for some $i$, $j$, then for all $t=0,1,\dots$}
@@
Let $p_{\min}^{t}$ and $p_{\max}^{t}$ be the minimum and maximum coordinates of the vector ${\bf{p}}^{t}$ respectively, and $L^{t}=\|{\bf{p}}^{t}\|^{2}$. Then for any $t=0,1,\dots$} the following inequalities hold:
@@
From Proposition 1 and the stochasticity of the vector ${\bf{p}}^{t}$, it follows that for any $t=0,1,\dots$
@@
The sequence $L^{t}=\|{\bf{p}}^{t}\|^{2}$, $t=0,1,\dots$ is decreasing.
@@
Consider the difference $L^{t+1}$ and $L^{t}$. Taking into account inequalities (6) we obtain
$L^{t+1}-L^{t}\leq p^{t+1}_{\max}-p^{t}_{\max}=\frac{p^{t}_{\max}(L^{t}-p^{t}_{\max})}{n-L^{t}}\leq 0.$
Thus, the sequence ${\{{L^{t}}\}}^{\infty}_{t=0}$ is decreasing. ∎
@@
The sequence $L^{t}=\|{\bf{p}}^{t}\|^{2}$, $t=0,1,\dots$ changes within the limits:
@@
$\mathcal{L}(p_{1},\dots,p_{n},\lambda)=\sum_{i=1}^{n}(p_{i})^{2}+\lambda\left(-1+\sum_{i=1}^{n}p_{i}\right),$
@@
Stationary points are found from the following system:
$\begin{cases}\frac{\partial\mathcal{L}}{\partial p_{1}}=2p_{1}+\lambda=0,\\ \frac{\partial\mathcal{L}}{\partial p_{2}}=2p_{2}+\lambda=0,\\ \cdots\\ \frac{\partial\mathcal{L}}{\partial p_{n}}=2p_{n}+\lambda=0,\\ \phi(p_{1},\dots,p_{n})=-1+\sum\limits_{i=1}^{n}p_{i}=0,\end{cases}\Longleftrightarrow\begin{cases}p_{1}=\frac{1}{n},\\ p_{2}=\frac{1}{n},\\ \cdots\\ p_{n}=\frac{1}{n},\\ \lambda=-\frac{2}{n}.\end{cases}$
Thus, we have a stationary point ${M^{\prime}}\big(\frac{1}{n},\frac{1}{n},\dots,\frac{1}{n},-\frac{2}{n}\big)$.
Let’s check the sufficient conditions for the extremum of the Lagrangian function. For this, we find the second partial derivatives of the function $\mathcal{L}(p_{1},\dots,p_{n},\lambda)$:
$\mathcal{L}^{{}^{\prime\prime}}_{p_{i}p_{i}}=2,\quad\mathcal{L}^{{}^{\prime\prime}}_{p_{i}p_{k}}=0,\quad i=\overline{1,n},\quad k=1,\ldots,n,\quad i\neq k.$
@@
${\rm d}^{2}\mathcal{L}\big(p_{1},\dots,p_{n},\lambda\big)=2{\rm d}{p_{1}^{2}}+2{\rm d}{p_{2}^{2}}+\cdots+2{\rm d}{p_{n}^{2}}=2\big({\rm d}{p_{1}^{2}}+{\rm d}{p_{2}^{2}}+\cdots+{\rm d}{p_{n}^{2}}\big).$
Since ${\rm d}^{2}\mathcal{L}>0$, the function $\mathcal{L}(p_{1},\dots,p_{n},\lambda)$ has a local minimum at point $M^{\prime}$}, $\mathcal{L}_{\min}=\mathcal{L}({M^{\prime}})=\frac{1}{n}$.
Thus, the function $L(p_{1},\dots,p_{n})$ under the constraint $\sum\limits_{i=1}^{n}p_{i}=1$ has a local conditional minimum at the point $M\left(\frac{1}{n},\frac{1}{n},\dots,\frac{1}{n}\right)$
$L_{\min}=L\left(\frac{1}{n},\frac{1}{n},\dots,\frac{1}{n}\right)=\frac{1}{n}.$
@@
The sequence $L^{t}=\|{\bf{p}}^{t}\|^{2}$, $t=0,1,\dots$ is convergent, i.e., there exists
@@
Figure 1: Trajectories of the dynamic conflict system for the case of three-dimensional space under the action of conflict transformation (4), i.e., the dynamics of change of arbitrary stochastic vectors ${\bf{p}}=(p_{1};p_{2};p_{3})$. The vertices of the triangle are identified with the coordinate sets $(1;0;0)$, $(0;1;0)$, and $(0;0;1)$. The limit values of the vectors $\bf{p}$} converge to the center of this triangle, meaning the center of the triangle is an attractor. The vertices of the triangle $P_{1}(1;0;0)$, $P_{2}(0;1;0)$, and $P_{3}(0;0;1)$ are unstable repellers. Only the center of the triangle, i.e., the point $\left(\frac{1}{3};\frac{1}{3};\frac{1}{3}\right)$, is stable. All other points of the triangle are unstable.
@@
$p_{1}^{\infty}=p_{2}^{\infty}=\dots=p_{j}^{\infty}=\dots=p_{n-1}^{\infty}=p_{n}^{\infty}.$ (8)
Therefore, the limit ${\bf{p}}^{\infty}=\lim\limits_{t\to\infty}{\bf{p}}^{t}$ exists. From the existence of the limit vector $\bf{p}$$^{\infty}$ and equations (4) it follows that ${\bf{r}}^{\infty}=\lim\limits_{t\to\infty}{\bf{r}}^{t}$ exists.
Suppose $\gamma(M)=0$. Considering (8) and the stochasticity of the vectors, we have for any $i=\overline{1,n}$:
@@
The calculation of the Jacobian matrix $J$ at the point ${\bf p}^{*}=\left(\frac{1}{n},\dots,\frac{1}{n}\right)$ yields a matrix with a very simple structure: all diagonal elements are identical ($J_{ii}=A$), and all off-diagonal elements are identical ($J_{ik}=B$). The Jacobian matrix $J$ has elements $J_{ik}=\frac{\partial p^{t+1}_{i}}{\partial p^{t}_{k}}$.
@@
Now we compute $A$ at the equilibrium point $\mathbf{p}^{*}=\left(\frac{1}{n},\dots,\frac{1}{n}\right)$. At this point, $p_{i}=\frac{1}{n}$. $L=\sum\limits_{j=1}^{n}\left(\frac{1}{n}\right)^{2}=n\cdot\left(\frac{1}{n^{2}}\right)=\frac{1}{n}$. Then $A=\frac{n^{3}-2n+2}{n(n^{2}-1)}.$
@@
Let’s compute $B$ at the equilibrium point ${\bf p}^{*}=\left(\frac{1}{n},\dots,\frac{1}{n}\right)$
@@
Consider the simplest boundary case where one component is zero, for example, $p_{n}=0$, and the remaining n1n-1 components are equal to $p_{i}^{*}=\frac{1}{n-1}$ for $i<n$.
For this point to be stable, it must be stable to perturbations that ”push” it off the boundary. That is, if we add a small amount $\varepsilon$ to the zero component $p_{n}$, it must decay rather than grow.
This stability is determined by the ”normal” eigenvalue, which is equal to the diagonal element of the Jacobian $J_{nn}$, calculated at this boundary point.
@@
Thus, stability analysis shows that only one limit point is stable: ${\bf{p}}^{\infty}={\bf{r}}^{\infty}=\left(\frac{1}{n},\frac{1}{n},\dots,\frac{1}{n}\right).$
@@
For the dynamic conflict system (1) with an initial stochastic vector ${\bf{p}}\in{\mathbb{R}}_{+}^{n}$}, $(n>1)$, whose evolution is given by formulas (5), the set of all possible limit vectors ${\bf p}^{\infty}\in\mathbb{R}_{+}^{n}$ has the following stability structure:
@@
    The fixed point ${\bf p}^{\infty}=\left(\frac{1}{n},\ldots,\frac{1}{n}\right)$ is a global attractor for all initial vectors ${\bf p}^{0}=\{p_{i}^{0}\}_{i=1}^{n}$}, where $p_{i}^{0}>0$ for all $i$.

### v2

# Mean–field dynamics of attractive resource interaction: From uniform to aggregated states

@@
Mean–field dynamics of attractive resource interaction: From uniform to aggregated states
We introduce and study a nonlinear discrete dynamical system describing the evolution of a resource distribution among interacting agents. The model generalizes several classical mean–field and opinion–dynamics frameworks and is defined on the standard simplex, where each coordinate evolves according to an interaction rule depending on preference–based mean–field interactions. We provide a complete analytical description of the long-term behavior of the system. First, we establish monotonicity properties and show that the dynamics always remains in a positively invariant region determined by initial conditions. We prove the existence of a unique fixed point for any admissible parameter set and derive an explicit closed-form formula for the equilibrium in arbitrary dimension. We then analyze the local stability of the fixed point and identify parameter regimes leading to aggregation or uniform distributions. Finally, we characterize all possible asymptotic scenarios and show that, despite the nonlinear structure, the system does not exhibit oscillatory or chaotic behavior: every trajectory converges to the unique equilibrium. The results provide a full qualitative theory for this class of monotone resource–interaction models and offer a mathematical explanation for the transition from uniform to aggregated states.
Keywords: discrete dynamical systems, attractive interaction, stability analysis, transcritical bifurcation, resource distribution, mean–field models.
@@ 1 Introduction
The type of interaction between components – repulsive or attractive – critically affects the system’s dynamics. Models with repulsive interaction, which often describe competition, typically lead to the system’s tendency to avoid resource concentration. In contrast, models with attractive interaction, which are the focus of this paper, can describe processes of cooperation, gravitation towards centers of influence, or capital concentration. It is such systems that are the subject of our investigation. While the term “conflict” might seem counterintuitive in a model with attractive interaction, it represents a competition for dominance or influence within the state space $\Omega$. In this context, attraction towards more influential components is a key strategic element of the conflict, leading to the aggregation of resources rather than their uniform dispersion. This connects to the foundational concept of “conflict” between probability measures introduced by Volodymyr Koshmanenko [4, 5]. While the original work focused on quantifying divergence, the attractive dynamics explored here demonstrate how such interactions can lead to “conflict resolution” through convergence towards a single, stable equilibrium. The mathematical modeling of interaction between conflicting systems, especially concerning the role of external support, is a subject of ongoing research [6].
@@
An alternative and powerful interpretation is as a system of interacting agents striving to reach consensus. In this view, the state vector represents a distribution of opinions, making the model an example of a discrete–time multi–agent system [10, 11]. Aggregated states then represent opinion polarization or fragmentation, phenomena widely studied in sociophysics [12, 13, 2].
Finally, the model shares properties with econophysics models of wealth distribution. The conservation law $\sum_{i}p_{i}^{t}=1$ mirrors a closed economy. While stochastic agent-based models like that of Chakraborti & Chakrabarti (2000) lead to Boltzmann–Gibbs distributions [14], our deterministic model provides an alternative mechanism for wealth aggregation, where the parameters $c_{i}$ can be interpreted as the “economic attractiveness” of an agent, analytically determining the final distribution.
Next, we construct a model of a dynamical system that describes the uniform or aggregated distribution of a certain resource in a given state space $\Omega=\{\omega_{1},\omega_{2},\ldots,\omega_{n}\}$, $n>1$. We fix a certain discrete probability measure $\mu$, which corresponds to a specific distribution. The value of this measure,
$\mu(\omega_{i})=:p_{i}\geq 0,\quad i=1,\ldots,n,$
is interpreted as the share of the resource or the relative weight of the i-th component of the system. By construction, $\sum_{i}p_{i}=1$. Therefore, the sequence $\mu(\omega_{i})$ defines a stochastic vector in ${\mathbb{R}}_{+}^{n}$: ${\bf p}=(p_{1},p_{2},\ldots,p_{n})$.
@@
It is easy to verify that these values also form a stochastic vector ${\bf r}=(r_{1},r_{2},\ldots,r_{n})\in{\mathbb{R}_{+}^{n}}$. The quantities $\nu(\omega_{i})$ can be interpreted as an average distribution (mean field) of $\mu$ over the subset $\Omega\setminus\omega_{i}$. This mean–field approach is widely used in statistical physics to describe systems where individual components interact with an average property of the entire system rather than with each other individually [15]. Furthermore, this paradigm has been substantially broadened by the theory of mean–field games, which provides a rigorous framework for analyzing strategic interactions within large populations of agents. Modern applications of this theory are highly diverse, ranging from ranking games with diffusion control [16] to spatial optimization on bounded domains [17]. While our model does not explicitly formulate a stochastic game, it shares the core mean–field philosophy, describing how the aggregate state of the system shapes the favorable conditions for individual components.
Conceptually, the interdependence of components through the global normalization constraint ($\sum_{i}p_{i}=1$) and the shared mean–field distribution establishes a profound link with the notion of correlated equilibrium. In classical game theory, a correlated equilibrium arises when players coordinate their strategies via a shared signal [18]. In the context of large populations and mean–field games, recent literature introduces mean–field analogues of correlated equilibrium, where the population distribution plays the role of a shared correlating signal [19, 20]. In our macroscopic system, the mean–field vector ${\bf r}^{t}$ and the normalizing denominator $z^{t}$ fulfill this exact function, naturally driving the system toward aggregated, cooperative outcomes rather than isolated competitive Nash equilibria.
@@
$p^{t+1}_{i}=\frac{p^{t}_{i}\big(1+{c_{i}}r^{t}_{i}\big)}{z^{t}},\quad 0<c_{i}\leq 1,\quad i=1,\ldots,n,$ (2)
where $r^{t}_{i}$ is the mean–field term defined as
$r^{t}_{i}=\frac{1-p^{t}_{i}}{n-1}.$ (3)
The parameters $c_{i}$ represent the time-invariant ($t$-independent) favorable conditions explicitly associated with each state $\omega_{i}$, thereby dictating the strength of the attractive interaction for that component. Finally, the normalizing denominator $z^{t}$, defined explicitly as
$z^{t}=\sum_{j=1}^{n}p^{t}_{j}\big(1+{c_{j}}r^{t}_{j}\big),$
ensures the stochasticity of the vector ${\bf p}^{t+1}$ (i.e., $\sum_{i=1}^{n}p_{i}^{t+1}=1$).
[…]
$p_{i}^{t+1}=\frac{1}{z^{t}}p_{i}^{t}(1-r_{i}^{t}),\quad r_{i}^{t+1}=\frac{1}{z^{t}}r_{i}^{t}(1-p_{i}^{t}),$
where $z^{t}=1-({\bf p}^{t},{\bf r}^{t})$ is the normalizing denominator (with $(\cdot,\cdot)$ denoting the scalar product in $\mathbb{R}^{n}$). The governing equation (2) studied in this work is a direct adaptation of this original equation, obtained by explicitly defining the opposing vector ${\bf r}^{t}$ as the mean–field distribution (3) and introducing the coefficients $c_{i}$ to scale the interaction strength.
@@
It can be noted that the “game for power” model proposed in [21] is a prime example of a system with repulsion, where agents seek to minimize the influence of others. The model presented in this manuscript, featuring attractive interaction ((2)), can thus be interpreted as a “game for cooperation” or alliance formation, where an agent’s benefit increases by gravitating towards resources controlled by others. The choice of attractive interaction fundamentally changes the system’s dynamics and significantly affects its properties.
@@
2 Uniformly distributed equilibria
Assume that in formulas (2), all $c_{i}=1$ for all $i=1,\ldots,n$, meaning favorable conditions are identical for each state $\omega_{i}$ and remain unchanged over time. Then equation (2) can be rewritten as
$p^{t+1}_{i}=\frac{p^{t}_{i}\big(1+r^{t}_{i}\big)}{1+\left({\bf p}^{t},{\bf r}^{t}\right)},\quad i=1,\ldots,n.$ (4)
@@ 2 Uniformly distributed equilibria
The fixed point where $p_{i}^{\infty}=\frac{1}{n}$ is known in the multi-agent systems literature as the consensus state [11, 23, 24]. It is a direct analogue to the outcome of classic linear consensus protocols, such as the DeGroot model, where iterative averaging on a fully connected graph leads to a system-wide average [25]. The model presented here, in this special case of uniform $c_{i}$, can thus be viewed as a novel, nonlinear consensus protocol. This perspective closely aligns with recent research, such as the work [26], which constructs a discrete-time opinion formation model of a similar type. In their framework, analogous to our uniform case, the iterative interaction mechanism effectively neutralizes competitive differences and drives the multi-agent system toward a state of absolute consensus.
@@
Every trajectory of the dynamic conflict system (1) with an arbitrary pair of initial stochastic vectors ${\bf p},{\bf r}\in{\mathbb{R}}_{+}^{n}$} $(n>1)$, whose evolution is given by formulas (5), converges to a fixed point
@@ 2. $2)$
The equilibrium is stable only in the case ${\bf p}^{\infty}={\bf r}^{\infty}=\left(\frac{1}{n},\frac{1}{n},\ldots,\frac{1}{n}\right)$.
@@
If $p_{i}\geq p_{j}$ for some $i$, $j$, then for all $t=0,1,\ldots$}
@@
Let $p_{\min}^{t}$ and $p_{\max}^{t}$ be the minimum and maximum coordinates of the vector ${\bf p}^{t}$ respectively, and $L^{t}=\|{\bf p}^{t}\|^{2}$. Then for any $t=0,1,\ldots$} the following inequalities hold:
@@
From Proposition 1 and the stochasticity of the vector ${\bf p}^{t}$, it follows that for any $t=0,1,\ldots$
@@
The sequence $L^{t}=\|{\bf p}^{t}\|^{2}$, $t=0,1,\ldots$ is decreasing.
@@
Let us express the difference $L^{t+1}-L^{t}$ using equations (5):
$L^{t+1}-L^{t}=\sum_{i=1}^{n}\left(p^{t}_{i}\cdot\frac{n-p^{t}_{i}}{n-L^{t}}\right)^{2}-L^{t}=\frac{1}{\big(n-L^{t}\big)^{2}}\sum_{i=1}^{n}\big(p^{t}_{i}\big)^{2}\left(\big(n-p^{t}_{i}\big)^{2}-\big(n-L^{t}\big)^{2}\right).$
Through algebraic rearrangement, $(n-p^{t}_{i})^{2}-(n-L^{t})^{2}=(L^{t}-p^{t}_{i})(2n-L^{t}-p^{t}_{i})$, and rewriting the second factor as $2(n-p^{t}_{i})-(L^{t}-p^{t}_{i})$, we expand the numerator:
$(n-L^{t})^{2}(L^{t+1}-L^{t})=2\sum_{i=1}^{n}p^{t}_{i}(L^{t}-p^{t}_{i})p^{t}_{i}(n-p^{t}_{i})-\sum_{i=1}^{n}\big(p^{t}_{i}\big)^{2}\big(L^{t}-p^{t}_{i}\big)^{2}.$
The second sum $\sum\limits_{i=1}^{n}\big(p^{t}_{i}\big)^{2}\big(L^{t}-p^{t}_{i}\big)^{2}\geq 0$, so its contribution to the difference is non-positive.
Considering the first sum, we define $q^{t}_{i}=p^{t}_{i}(L^{t}-p^{t}_{i})$ and $f(p_{i}^{t})=p_{i}^{t}(n-p_{i}^{t})$. Notice that the sum of the elements $q^{t}_{i}$ is exactly zero:
$\sum_{i=1}^{n}q^{t}_{i}=L^{t}\sum_{i=1}^{n}p^{t}_{i}-\sum_{i=1}^{n}\big(p^{t}_{i}\big)^{2}=L^{t}\cdot 1-L^{t}=0.$
Since $n\geq 2$ and $p_{i}^{t}\in[0,1]$, the derivative $f^{\prime}(p_{i}^{t})=n-2p_{i}^{t}\geq 0$, meaning $f(p_{i}^{t})$ is strictly increasing on $[0,1]$.
Let us evaluate the product $q^{t}_{i}f(p^{t}_{i})$ relative to the constant $f(L^{t})$. Since $f$ is increasing on $[0,1]$, we consider two possible cases.
Case 1: $p^{t}_{i}\leq L^{t}$. In this case $q^{t}_{i}=p^{t}_{i}(L^{t}-p^{t}_{i})\geq 0$ and, by the monotonicity of $f$,
$f(p^{t}_{i})\leq f(L^{t}).$
Therefore,
$q^{t}_{i}f(p^{t}_{i})\leq q^{t}_{i}f(L^{t}).$
Case 2: $p^{t}_{i}>L^{t}$. Then $q^{t}_{i}<0$ and $f(p^{t}_{i})>f(L^{t})$. Multiplying this inequality by the negative factor $q^{t}_{i}$ reverses its direction, which yields
$q^{t}_{i}f(p^{t}_{i})<q^{t}_{i}f(L^{t}).$
Thus, in both cases the inequality
$q^{t}_{i}f(p^{t}_{i})\leq q^{t}_{i}f(L^{t})$
holds for all $i=1,\ldots,n$. Summing over $i$ gives
$\sum_{i=1}^{n}q^{t}_{i}f(p^{t}_{i})\leq\sum_{i=1}^{n}q^{t}_{i}f(L^{t})=f(L^{t})\sum_{i=1}^{n}q^{t}_{i}=f(L^{t})\cdot 0=0.$
Since both terms in the expansion of $L^{t+1}-L^{t}$ are non-positive, we conclude that
$L^{t+1}-L^{t}\leq 0.$
Thus, the sequence $\{L^{t}\}_{t=0}^{\infty}$ is decreasing. ∎
@@
The sequence $L^{t}=\|{\bf p}^{t}\|^{2}$, $t=0,1,\ldots$ changes within the limits:
@@
$\mathcal{L}(p_{1},\ldots,p_{n},\lambda)=\sum_{i=1}^{n}(p_{i})^{2}+\lambda\left(-1+\sum_{i=1}^{n}p_{i}\right),$
@@
Critical points are found from the following system:
$\begin{cases}\frac{\partial\mathcal{L}}{\partial p_{1}}=2p_{1}+\lambda=0,\\ \frac{\partial\mathcal{L}}{\partial p_{2}}=2p_{2}+\lambda=0,\\ \cdots\\ \frac{\partial\mathcal{L}}{\partial p_{n}}=2p_{n}+\lambda=0,\\ \phi(p_{1},\ldots,p_{n})=-1+\sum\limits_{i=1}^{n}p_{i}=0,\end{cases}\Longleftrightarrow\begin{cases}p_{1}=\frac{1}{n},\\ p_{2}=\frac{1}{n},\\ \cdots\\ p_{n}=\frac{1}{n},\\ \lambda=-\frac{2}{n}.\end{cases}$
Thus, we have a fixed point ${M^{\prime}}\big(\frac{1}{n},\frac{1}{n},\ldots,\frac{1}{n},-\frac{2}{n}\big)$.
Let’s check the sufficient conditions for the extremum of the Lagrangian function. For this, we find the second partial derivatives of the function $\mathcal{L}(p_{1},\ldots,p_{n},\lambda)$:
$\mathcal{L}^{{}^{\prime\prime}}_{p_{i}p_{i}}=2,\quad\mathcal{L}^{{}^{\prime\prime}}_{p_{i}p_{k}}=0,\quad i=1,\ldots,n,\quad k=1,\ldots,n,\quad i\neq k.$
@@
${\rm d}^{2}\mathcal{L}\big(p_{1},\ldots,p_{n},\lambda\big)=2{\rm d}{p_{1}^{2}}+2{\rm d}{p_{2}^{2}}+\cdots+2{\rm d}{p_{n}^{2}}=2\big({\rm d}{p_{1}^{2}}+{\rm d}{p_{2}^{2}}+\cdots+{\rm d}{p_{n}^{2}}\big).$
Since ${\rm d}^{2}\mathcal{L}>0$, the function $\mathcal{L}(p_{1},\ldots,p_{n},\lambda)$ has a local minimum at point $M^{\prime}$}, $\mathcal{L}_{\min}=\mathcal{L}({M^{\prime}})=\frac{1}{n}$.
Thus, the function $L(p_{1},\ldots,p_{n})$ under the constraint $\sum\limits_{i=1}^{n}p_{i}=1$ has a local conditional minimum at the point $M\left(\frac{1}{n},\frac{1}{n},\ldots,\frac{1}{n}\right)$
$L_{\min}=L\left(\frac{1}{n},\frac{1}{n},\ldots,\frac{1}{n}\right)=\frac{1}{n}.$
@@
The sequence $L^{t}=\|{\bf p}^{t}\|^{2}$, $t=0,1,\ldots$ is convergent, i.e., there exists
@@
Figure 1: Phase portraits on the 2D invariant simplex within the 3D state space, illustrating the asymptotic trajectories of the stochastic vectors ${\bf p}^{t}$ under the action of conflict transformation (4). The trajectories asymptotically converge to the barycenter of the simplex $\left(\frac{1}{3},\frac{1}{3},\frac{1}{3}\right)$, which constitutes the interior attractor. The vertices of the simplex $P_{1}(1;0;0)$, $P_{2}(0;1;0)$, and $P_{3}(0;0;1)$ are unstable repelling fixed points. Saddle points $S_{12}(0.5;0.5;0)$, $S_{23}(0;0.5;0.5)$, and $S_{13}(0.5;0;0.5)$ are located at the midpoints of the one-dimensional faces (edges) of the simplex. Their stable invariant manifolds coincide strictly with the boundary edges, capturing the phase flow from the repelling vertices. Concurrently, their unstable invariant manifolds extend into the interior of the phase space, defining the principal pathways of the flow toward the central attractor. All other points within the simplex represent transient states.
@@
$p_{1}^{\infty}=p_{2}^{\infty}=\ldots=p_{j}^{\infty}=\ldots=p_{n-1}^{\infty}=p_{n}^{\infty}.$ (8)
Therefore, the limit ${\bf p}^{\infty}=\lim\limits_{t\to\infty}{\bf p}^{t}$ exists. From the existence of the equilibrium vector $\bf p$$^{\infty}$ and equations (4) it follows that ${\bf r}^{\infty}=\lim\limits_{t\to\infty}{\bf r}^{t}$ exists.
Suppose $\gamma(M)=0$. Considering (8) and the stochasticity of the vectors, we have for any $i=1,\ldots,n$:
@@
The calculation of the Jacobian matrix $J$ at the point ${\bf p}^{*}=\left(\frac{1}{n},\ldots,\frac{1}{n}\right)$ yields a matrix with a very simple structure: all diagonal elements are identical ($J_{ii}=A$), and all off-diagonal elements are identical ($J_{ik}=B$). The Jacobian matrix $J$ has elements $J_{ik}=\frac{\partial p^{t+1}_{i}}{\partial p^{t}_{k}}$}.
@@
Now we compute $A$ at the equilibrium point $\mathbf{p}^{*}=\left(\frac{1}{n},\ldots,\frac{1}{n}\right)$. At this point, $p_{i}=\frac{1}{n}$. $L=\sum\limits_{j=1}^{n}\left(\frac{1}{n}\right)^{2}=n\cdot\left(\frac{1}{n^{2}}\right)=\frac{1}{n}$. Then $A=\frac{n^{3}-2n+2}{n(n^{2}-1)}.$
@@
Let’s compute $B$ at the equilibrium point ${\bf p}^{*}=\left(\frac{1}{n},\ldots,\frac{1}{n}\right)$
@@
Consider the simplest boundary case where one coordinate is zero, for example, $p_{n}=0$, and the remaining n1n-1 coordinates are equal to $p_{i}^{*}=\frac{1}{n-1}$ for $i<n$.
For this point to be stable, it must be stable to perturbations that “push” it off the boundary. That is, if we add a small positive perturbation $\varepsilon$ to the zero coordinate $p_{n}$, it must decay rather than grow.
This stability is determined by the “normal” eigenvalue, which is equal to the diagonal element of the Jacobian $J_{nn}$, calculated at this boundary point.
@@
Thus, stability analysis shows that only one equilibrium is stable: ${\bf p}^{\infty}={\bf r}^{\infty}=\left(\frac{1}{n},\frac{1}{n},\ldots,\frac{1}{n}\right).$
@@
For the dynamic conflict system (1) with an initial stochastic vector ${\bf p}\in{\mathbb{R}}_{+}^{n}$}, $(n>1)$, whose evolution is given by formulas (5), the set of all possible vectors of equilibrium ${\bf p}^{\infty}\in\mathbb{R}_{+}^{n}$ has the following stability structure:
@@
    The interior fixed point ${\bf p}^{\infty}=\left(\frac{1}{n},\ldots,\frac{1}{n}\right)$ is a global attractor for all initial vectors ${\bf p}^{0}=\{p_{i}^{0}\}_{i=1}^{n}$}, where $p_{i}^{0}>0$ for all $i$.